OpenStudy (anonymous):

george and carmen went on a bicycle trip. they took a bus to their starting point and then biked the rest. they traveled 350 kilometers in total and they biked 80 kilometers more than they bussed let x= kilometer traveled by biked and y= kilometer traveled by bus Find how many kilometer the traveled by bike?

OpenStudy (anonymous):

help me please i dont understand

OpenStudy (anonymous):

x = bike y = bus x + y = 350 x = y + 80 y + 80 + y = 350 2y = 350 - 80 2y = 270 y = 135 x + y = 350 x + 135 = 350 x = 350 - 135 x = 215 check.. x = y + 80 215 = 135 + 80 215 = 215 (correct) check... x + y = 350 215 + 135 = 350 350 = 350 (correct) They rode their bikes 215 km,and they rode the bus 135 km. Any questions ?

OpenStudy (anonymous):

thanks i understand now